OpenStudy (anonymous):

Determine the solution set of (2x - 1)2 - 100 = 0

OpenStudy (llort):

I am getting x=25.5

OpenStudy (llort):

((51*2)/2-1)2-100=0

OpenStudy (anonymous):

4x-2-100=0 4x-102=0 4x=102 x=102/4

OpenStudy (a_clan):

\[(2x-1)^{2} - 100 = 0\] \[(2x-1)^{2} = 100\] So \[(2x-1) = +10 \] \[or(2x-1) = -10\] So x=11/2, -9/2

jhonyy9 (jhonyy9):

2x=51 x=51/2 x=25,5